OpenStudy (anonymous):

How do we " find p (rolling 2 or 5) with one number cube?"

OpenStudy (ganpat):

cube has 6 sides.. so chances of getting 2 = 1/6 as you have only one 2 similarly, getting 5 =1/6 so, (1/6) * (1/6) = 1/36..

OpenStudy (anonymous):

@ganpat the question is 2 OR 5, not 2 then a 5, the chances are 2/6, reduce that to 1/3

OpenStudy (ganpat):

oh sorry, it should be (1/6) + (1/6) = 2/6 = 1/3...